# Magnetic Skyrmions: from lumps to supercompactons

### v1

@@ Abstract
The magnetic Skyrmion is described by one control parameter and one length scale. We study the two extreme limits of the control parameter – infinitely large and vanishing – and find that the magnetic Skyrmion becomes a “restricted” magnetic Skyrmion and an $\operatorname{O}(3)$ sigma model lump, respectively. Depending on the potential under consideration, the restricted limit manifests differently. In the case of the Zeeman term, the restricted magnetic Skyrmion becomes a “supercompacton” that develops a discontinuity, whereas for the Zeeman term to the power 3/2 it becomes a normal compacton. Finally, we observe that the case of the Zeeman term squared, which can also be understood as a special combination of the Zeeman term and the easy-plane potential – realizable in the laboratory, the lump solution is an analytically exact solution for all values of the control parameter.
@@ 1 Introduction
In this work we consider the theory of magnetic Skyrmions in two spatial dimension (2D) and apply certain techniques used in the past in the context of the baby-Skyrme model [42]. We show that a dimensionless parameter enables us to control the relation between the Dirichlet term, the one with two derivatives, and the rest of the energy functional that contains the potential and the DM term. By changing this parameter we can flow from an almost BPS theory, close to the $\operatorname{O}(3)$ sigma model which admits lumps as solutions, to a “restricted” magnetic Skyrmion model with just $E_{1}$ and $E_{0}$. The restricted model is analytically solvable. In the case of the ordinary Zeeman potential, which is quadratic in the perturbative excitation fields near the vacuum, a feature never encountered before in the context of the baby-Skyrme model is present: the solution is not only a compacton but it is also discontinuous. We may call it a “supercompacton”. The discontinuity is resolved as soon as we turn on a small coefficient for the Dirichlet term. We also study another subtlety of the model; that is, the DM term can be integrated by parts without changing the equation of motion or the energy when the potential is the Zeeman term (giving rise to exponential falloff). However, it turns out that for potentials that give rise to massless perturbative excitations, e.g. the Zeeman squared term, which can be understood as a specific linear combination of the Zeeman term and the easy-plane potential, discarding a boundary term and using a different DM term in the energy functional, is crucial for computing the correct sizes of the solutions in the lump limit, i.e. the limit of the control parameter going to zero.
@@
Finally, we observe that this lump-type solution corresponds to the solution of the critically coupled version of magnetic Skyrmions with the Zeeman term squared, as described by Barton-Singer et al[19]. However, it is not limited to this scenario: it is also the solution to the full second-order variational equation of motion for all values of the coupling, albeit with a different “lump size” that we determine. This exact solution is thus a solution in the case of the Zeeman term squared for all values of the coupling and hence may be advantageous for further studies.
@@
On the experimental side, we also confront the dimensionless parameter with the various experimental realizations of the magnetic Skyrmion and find that the (near) “restricted” magnetic Skyrmion – corresponding to the case of a large control parameter – is actually realized in some materials. We also confirm that it is possible to experimentally realize the Zeeman squared term, by using magnets with a certain anisotropy term and adjusting the magnetic field of the Zeeman term, so as to complete the square. At this critically tuned value of the magnetic field, the exact analytic lump-type solution is thus realizable in the laboratory.
@@ 2.3 $\operatorname{O}(3)$ sigma model limit
We notice that the correction to the energy is linearly proportional to $\epsilon$.
@@ 2.4 Restricted model limit
In the case of $\ell$ fixed and $\epsilon\to\infty$ we expect that the stable Skyrmion is essentially given by a stabilization between $E_{1}$ and $E_{0}$, while $E_{2}$ is negligible with respect to the other two terms. Thus the energy functional takes the simplified form111 One may wonder why we restrict the potential to the form $V_{p}$ of Eq. (2.5) and hence do not consider e.g. $V\propto(1-(n^{3})^{2})$. Although such easy-axis potential is often utilized in magnetic systems, they do not allow for the large-$\epsilon$ limit because the ground state becomes inhomogeneous [45, 46] and requires a different analysis than carried out here.
@@
One interesting aspect of these equations is that they are first order in the derivative. Thus, they can easily be solved analytically. Unlike the restricted baby-Skyrme model [47, 48, 42], here we do not have an infinite dimensional symmetry group. The restricted equation for the hedgehog Ansatz has no derivatives and can be written as
@@ 2.6 Analytic solution at any coupling
We note that the Bogomol’nyi completion requires us to select the critical coupling $h=\frac{\kappa^{2}}{2}$ enabling us to complete the square in the energy functional. This choice of coupling corresponds to $\epsilon=2$. However, another interesting fact about the Zeeman squared potential ($p=2$) is that inserting the Ansatz $\vartheta=2\arctan\tfrac{a}{r}$ into the full second-order Euler-Lagrange equation gives
@@
which beautifully has the solution $a=\frac{\kappa}{h}=\ell$. This size does not contradict the BPS solution, since the first-order BPS equation required $h=\frac{\kappa^{2}}{2}$ which in the full second-order solution corresponds to the lump size $a=\frac{\kappa}{h}=\frac{2}{\kappa}$. It is interesting, however to insert the exact solution into the energy functional, which gives
@@
The $p=2$ case is thus applicable to many more physical systems, than just the critically coupled one. Having the analytic solution in hand may be advantageous for further studies, including interaction and scattering etc.
@@ 2.7 Numerical solutions
Figure 5: Energy density for: (a) $p=\frac{3}{2}$ (b) $p=1$ and $\epsilon=0.001,0.01,0.1,1$, and in (c) $p=\frac{3}{2}$ (d) $p=1$ and $\epsilon=1,10,100,1000$.
@@
Fig. 5 shows the energy densities for the sigma-model lump limit ($\epsilon\to 0$) in panels (a) and (b) for $p=\frac{3}{2}$ and $p=1$, respectively, and for the restricted limit ($\epsilon\to\infty$) in panels (c) and (d) for $p=\frac{3}{2}$ and $p=1$, respectively. In both the charge density as well as in the energy density, the discontinuity that develops in the $p=1$ case for $\epsilon\to\infty$ is clearly seen in Figs. 4(b) and 5(d), respectively.
@@
Figure 6: Energy in the lump limit ($\epsilon\to 0$) as a function of $\epsilon$ for $p=\tfrac{3}{2}$ and $p=1$. The blue dashed line corresponds to the analytical energy formula (2.46) for the lump in the limit of small but nonvanishing $\epsilon$ in the case of $p=\frac{3}{2}$. The red dashed line instead is the energy of Eq. (2.51) with the lump size taken from Fig. 3.
@@
We show the total energy of the numerical solutions in the lump limit ($\epsilon\to 0$) for the $p=\frac{3}{2},1$ cases in Fig. 6. We notice that the $p=\frac{3}{2}$ case is well described by the energy functional (2.46) with the correct lump size (2.45). This confirms that discarding the boundary term in Eq. (2.40) is appropriate for computing the energies in the lump limit. In the $p=1$ case, we show the regularized energy (2.47) with a dashed red curve and lump size $a$ as chosen in Fig. 3.
Figure 7: (a) The profile function $\vartheta$ divided by the lump solution $\vartheta_{\rm lump}$ of Eq. (2.43) for $\epsilon=10^{-k}$ with $k=4,\ldots,8$. (b) The lump size $c=\frac{a}{\ell}$ estimated from the derivative $\vartheta^{\prime}(0)=-\frac{2}{c\ell}$.
Finally, we consider the difficult problem of estimating the lump size in the $\epsilon\to 0$ limit. Since we can see from Fig. 3 that the profiles match the lump solution if the lump size is fitted correctly, we can extract this lump size from the numerical solutions by taking the derivative of the solution at $r=0$. This yields $\vartheta^{\prime}(0)=-\frac{2}{a}=-\alpha$, where the latter is the shooting parameter and this parameter is known quite accurately from the numerical solution. In Fig. 7(a) the ratio of the numerical solutions $\vartheta$ and the exact lump solution (2.43) is shown to tend to unity for $\epsilon\to 0$. The smallest $\epsilon$ in the figure is $\epsilon=10^{-8}$. Fig. 7(b) instead shows the numerically extracted lump sizes and a power-law fit is made for the smallest-$\epsilon$ part of the figure, finding a power-law for the lump size as $a\approx 4.7\ell\epsilon^{0.29}$. Due to the difficulty of very small couplings, it is hard to tell whether the lump size asymptotically would tend to a power-law like $\epsilon^{1/3}$ or $\epsilon^{1/4}$ or something else.
@@ 2.8 Comparison with experiments
where $\operatorname{J}$ is the unit Joule, not to be confused with the interaction strength of the Heisenberg Hamiltonian. Note that in general $\ell$ can change significantly, viz. one order of magnitude. For instance, if the we fix $\ell=100\operatorname{nm}$ and $\lambda\in\mathcal{O}(10^{1})$ nm as before, we find that $\epsilon$ changes one order of magnitude, i.e. $\epsilon\simeq 100$. For materials other than MnSi, $\lambda$ can also be of the order $\mathcal{O}(10^{2})\operatorname{nm}$. Therefore, considering the other two cases where $\lambda$ is fixed as $\lambda\sim 10^{2}\operatorname{nm}$ and $\ell$ takes its maximum and minimum values, we find the relevant ranges for $\epsilon$:
@@
In Secs. 2.5 and 2.6, we demonstrated that in the special physical case of $p=2$, which describes chiral magnets with anisotropy interactions, any spherical magnetic Skyrmion can be represented by the analytical lump solution, regardless of its DM coupling, $\kappa$. However, the potential $V_{p=2}$ corresponds to the sum of a magnetic field and anisotropy (easy-plane) contributions. To realize this potential, the couplings of these terms must be equal. It turns out to be possible [50]. Furthermore, since the critical coupling regime corresponds to the condition $\forall\leavevmode\nobreak\ \ell,\ \epsilon=2$, which satisfies the inequalities (2.72), we suggest that, in principle, a BPS magnetic Skyrmion could be realized experimentally.
@@ 3 Conclusion
In this paper, we have studied magnetic Skyrmions in chiral magnets with three different potentials, corresponding to $p=1$, $p=\frac{3}{2}$ and $p=2$ in the limits of the control parameter $\epsilon=\frac{\kappa^{2}}{h}=\frac{4D^{2}}{|B|J}$ becoming very large and going to zero. The $p=1$ case is simply the Zeeman potential term and the $p=2$ case is a particular combination of the easy-plane (anisotropy) and the Zeeman potentials. The $p=\frac{3}{2}$ case is probably somewhat more academic. The $\epsilon$ vanishing limit reduces the Skyrmion to a sigma model lump with a lump size that we accurately estimated, except for the case of the Zeeman term ($p=1$). The large-$\epsilon$ limit instead corresponds to the “restricted” magnetic Skyrmions, for which we found the exact analytic solutions as well. In the $p=\frac{3}{2}$ case, the restricted solution is a compacton, similar in nature to those found in the baby-Skyrme [48] and 3D Skyrme models [51]. However, in the $p=1$ case corresponding to the Zeeman term, we find a new type of restricted solution, which we dub a supercompacton, since it becomes discontinuous. We provide numerical solutions showing that they slowly tend to the supercompacton limit for very large values of $\epsilon$. We have also reviewed the beautiful result of the critically coupled magnetic Skyrmion of Barton-Singer-Ross-Schroers [19], where an analytic solution is available, hence providing many studies with a clear advantage. Their solution is limited to the critical coupling. However, in this paper, we demonstrate that the same solution is applicable to any value of the coupling parameter, not solely the critical one, providing an exact analytic solution for all systems incorporating the Zeeman squared term.

### v2

@@ Abstract
The magnetic Skyrmion is described by one control parameter and one length scale. We study the two extreme limits of the control parameter – infinitely large and vanishing – and find that the magnetic Skyrmion becomes a “restricted” magnetic Skyrmion and an $\operatorname{O}(3)$ sigma model lump, respectively. Depending on the potential under consideration, the restricted limit manifests differently. In the case of the Zeeman term, the restricted magnetic Skyrmion becomes a “supercompacton” that develops a discontinuity, whereas for the Zeeman term to the power 3/2 it becomes a normal compacton. In both the lump and the restricted limit the solution is given in exact explicit form. We observe that the case of the Zeeman term squared, which can also be understood as a special combination of the Zeeman term and the easy-plane potential – realizable in the laboratory, the analytically exact solution for all values of the coupling – including the BPS case – is also of the lump type. Finally, we notice that certain materials (e.g. Fe1–xCoxSi or Mn1–xFexGe) have a rather large control parameter $\epsilon$ of order 100, making the restricted limit a suitable rough approximation.
@@ 1 Introduction
In this work we consider the theory of magnetic Skyrmions in two spatial dimensions (2D) and apply certain techniques used in the past in the context of the baby-Skyrme model [42]. We show that a dimensionless parameter enables us to control the relation between the Dirichlet term, the one with two derivatives, and the rest of the energy functional that contains the potential and the DM term. By changing this parameter we can flow from an almost BPS theory, close to the $\operatorname{O}(3)$ sigma model which admits lumps as solutions, to a “restricted” magnetic Skyrmion model with just $E_{1}$ and $E_{0}$. The restricted model is analytically solvable. In the case of the ordinary Zeeman potential, which is quadratic in the perturbative excitation fields near the vacuum, a feature never encountered before in the context of the baby-Skyrme model is present: the solution is not only a compacton111 Compactons are solitonic solutions with compact support: they have a nontrivial profile function for $r<R$ and vanishing energy outside, i.e. for $r>R$. Often the fields are continuous but their derivatives are not at $r=R$. Compactons exist in the baby-Skyrme model [43, 44, 45], the BPS-Skyrme model [46], as Q-balls in the gauged (ungauged) $\mathbb{C}P^{N}$ model [47]([48]) and in K-field theories [49], etc. but it is also discontinuous. We may call it a “supercompacton”. The discontinuity is resolved as soon as we turn on a small coefficient for the Dirichlet term. We also study another subtlety of the model; that is, the DM term can be integrated by parts without changing the equation of motion or the energy when the potential is the Zeeman term (giving rise to exponential falloff). However, it turns out that for potentials that give rise to massless perturbative excitations, e.g. the Zeeman squared term, which can be understood as a specific linear combination of the Zeeman term and the easy-plane potential, discarding a boundary term and using a different DM term in the energy functional, is crucial for computing the correct sizes of the solutions in the lump limit, i.e. the limit of the control parameter going to zero.
@@
Finally, we observe that this lump-type solution corresponds to the solution of the critically coupled version of magnetic Skyrmions with the Zeeman term squared, as described by Barton-Singer et al[19]. It is not limited to this scenario: it is also the solution to the full second-order variational equation of motion for all values of the coupling [19]. We find that for this potential, discarding the boundary term changes the energy functional to become dependent on $\epsilon$, hence revealing a phase transition between the homogeneous phase and the Skyrmionic phase.
@@
On the experimental side, we also confront the dimensionless parameter with the various experimental realizations of the magnetic Skyrmion and find that the (near) “restricted” magnetic Skyrmion – corresponding to the case of a large control parameter – is actually realized in some materials, like Fe1–xCoxSi or Mn1–xFexGe. We also confirm that it is possible to experimentally realize the Zeeman squared term, by using magnets with a certain anisotropy term and adjusting the magnetic field of the Zeeman term, so as to complete the square. At this critically tuned value of the magnetic field, the exact analytic lump-type solution is thus realizable in the laboratory.
@@ 2.3 $\operatorname{O}(3)$ sigma model limit
[…]
@@ 2.4 Restricted model limit
In the case of $\ell$ fixed and $\epsilon\to\infty$ we expect that the stable Skyrmion is essentially given by a stabilization between $E_{1}$ and $E_{0}$, while $E_{2}$ is negligible with respect to the other two terms. Thus the energy functional takes the simplified form333 One may wonder why we restrict the potential to the form $V_{p}$ of Eq. (2.5) and hence do not consider e.g. $V\propto(1-(n^{3})^{2})$. Although such easy-axis potential is often utilized in magnetic systems, they do not allow for the large-$\epsilon$ limit because the ground state becomes inhomogeneous [52, 53] and requires a different analysis than carried out here.
@@
One interesting aspect of these equations is that they are first order in the derivative. Thus, they can easily be solved analytically. Unlike the restricted baby-Skyrme model [43, 44, 45, 42], here we do not have an infinite dimensional symmetry group. The restricted equation for the hedgehog Ansatz has no derivatives and can be written as
@@ 2.5 Critical coupling
We now consider what happens if we again drop the boundary term of Eq. (2.16) as we had to in the lump case of Sec. 2.3. The energy is thus changed by a total derivative, yielding now
$E$ $=2\pi\int r\mathop{}\!\mathrm{d}r\left[\frac{1}{2}\left(\vartheta^{\prime}-\frac{\sin\vartheta}{r}+\kappa(1-\cos\vartheta)\right)^{2}+\frac{\sin\vartheta}{r}\vartheta^{\prime}+\frac{\kappa}{r}\partial_{r}(r\sin\vartheta(1-\cos\vartheta))\right]$
$=\pi\int r\mathop{}\!\mathrm{d}r\left(\vartheta^{\prime}-\frac{\sin\vartheta}{r}+\kappa(1-\cos\vartheta)\right)^{2}+4\pi Q+4\pi\Omega^{\prime}\ .$ (2.67)
Notice that a total derivative other than the topological charge remains
$\Omega^{\prime}=\frac{\kappa}{2}\int\mathop{}\!\mathrm{d}r\,\partial_{r}(r\sin\vartheta(1-\cos\vartheta))=0\ .$ (2.68)
which however vanishes when evaluated on the BPS solution. This means that dropping the boundary term in Eq. (2.16) leads to a Bogomol’nyi energy (bound)
$E\geq-4\pi\ .$ (2.69)
@@ 2.6 Analytic solution at any coupling
We note that the Bogomol’nyi completion requires us to select the critical coupling $h=\frac{\kappa^{2}}{2}$ enabling us to complete the square in the energy functional. This choice of coupling corresponds to $\epsilon=2$. However, another interesting fact about the Zeeman squared potential ($p=2$) is that inserting the Ansatz $\vartheta=2\arctan\tfrac{a}{r}$ into the full second-order Euler-Lagrange equation gives [19]
@@
which beautifully has the solution $a=\frac{\kappa}{h}=\ell$. This size does not contradict the BPS solution, since the first-order BPS equation required $h=\frac{\kappa^{2}}{2}$ which in the full second-order solution corresponds to the lump size $a=\frac{\kappa}{h}=\frac{2}{\kappa}$. It is interesting, however to insert the exact solution into the energy functional (2.35), which gives
@@
The $p=2$ case is thus applicable to many more physical systems, than just the critically coupled one [19]. Having the analytic solution in hand may be advantageous for further studies, including interaction and scattering etc.
In Sec. (2.3), we found however that the lump-type Ansatz gave the incorrect lump size $a$ by using the energy (2.35), and we should instead drop the boundary term of Eq. (2.40) yielding the energy
$E=2\pi\int r\mathop{}\!\mathrm{d}r\;\frac{4h\kappa^{2}(h^{3}r^{2}+h(1-3hr^{2})\kappa^{2}+\kappa^{4})}{(h^{2}r^{2}+\kappa^{2})^{3}}=4\pi(1-\epsilon)\ .$ (2.72)
We may interpret the change of sign of the energy at $\epsilon=\epsilon^{\rm crit}=1$ as the Skyrmionic phase being energetically preferable for $\epsilon>1$ and the homogeneous phase on the other hand preferable for $\epsilon<1$. Luckily, the BPS-case resides in the Skyrmionic phase at $\epsilon=2$.
@@ 2.7 Numerical solutions
(e) $p=\tfrac{3}{2}$
(f) $p=1$
Figure 5: Energy density for: (a) $p=\frac{3}{2}$ (b) $p=1$ and $\epsilon=0.001,0.01,0.1,1$ with the boundary term of Eq. (2.40) (solid blue lines) and without it (dashed dark-blue lines), and in (c) $p=\frac{3}{2}$ (d) $p=1$ and $\epsilon=1,10,100,1000$ with the boundary term and in (e) $p=\frac{3}{2}$ (f) $p=1$ and $\epsilon=1,10,100,1000$ without the boundary term of Eq. (2.40).
@@
Fig. 5 shows the energy densities for the sigma-model lump limit ($\epsilon\to 0$) in panels (a) and (b) with and without the boundary term for $p=\frac{3}{2}$ and $p=1$, respectively, and for the restricted limit ($\epsilon\to\infty$) in panels (c) and (d) for $p=\frac{3}{2}$ and $p=1$, respectively, with the boundary term and in panels (e) and (f) $p=\frac{3}{2}$ and $p=1$, respectively, without the boundary term of Eq. (2.40). In both the charge density as well as in the energy density, the discontinuity that develops in the $p=1$ case for $\epsilon\to\infty$ is clearly seen in Figs. 4(b) and 5(d), respectively. We note that the energy density is divergent (singular) in the $p=\frac{3}{2}$ case at $\rho=1$ when the boundary term is taken into account, see Fig. 5(c), but is convergent (finite) when the boundary term is dropped, see Fig. 5(e). We also note that the total energy (the integral) is unchanged in the $p=1$ case (see Figs. 5(d) and 5(f)), but the energy density is slightly modified by the presence of the boundary term.
@@
Figure 6: Energy in the lump limit ($\epsilon\to 0$) as a function of $\epsilon$ for $p=\tfrac{3}{2}$ and $p=1$. (a) The blue dashed line corresponds to the analytical energy formula (2.46) for the lump in the limit of small but nonvanishing $\epsilon$ in the case of $p=\frac{3}{2}$. The red dashed line instead is the energy of Eq. (2.51) with the lump size taken from Fig. 3. (b) The energy shown up to larger values of $\epsilon$ revealing the sign change near $21<\epsilon<$ 2. The case $p=\frac{3}{2}$ is shown with (solid blue line) and without the boundary term (dashed blue line) of Eq. (2.40).
@@
We show the total energy of the numerical solutions in the lump limit ($\epsilon\to 0$) for the $p=\frac{3}{2},1$ cases in Fig. 6(a). We notice that the $p=\frac{3}{2}$ case is well described by the energy functional (2.46) with the correct lump size (2.45). This confirms that discarding the boundary term in Eq. (2.40) is appropriate for computing the energies in the lump limit. In the $p=1$ case, we show the regularized energy (2.47) with a dashed red curve and lump size $a$ as chosen in Fig. 3. We show the sign change of the energy in the $p=\frac{3}{2},1$ case in Fig. 6(b) which appears to take place around $21<\epsilon<$ 2, which is consistent with the lump prediction (2.46) which for $p=\frac{3}{2}$ should take place at $\epsilon=\epsilon^{\rm crit}=\sqrt{2}$.
Figure 7: The lump size $c=\frac{a}{\ell}$ estimated by fitting the lump solution to the numerical solutions. The dashed black curve is a logarithmic fit.
Finally, we consider the difficult problem of estimating the lump size in the $\epsilon\to 0$ limit. Since we can see from Fig. 3 that the profiles match the lump solution if the lump size is fitted correctly, we can extract this lump size from the numerical solutions by fitting the size of the lump solution to the numerical ones. Fig. 7 shows the numerically extracted lump sizes and a logarithmic fit finding an approximate behavior of the lump size as $c\approx 1.291(-\log\epsilon)^{-0.918}$. Clearly the lump size goes very slowly to zero.
@@ 2.8 Comparison with experiments
where $\operatorname{J}$ is the unit Joule, not to be confused with the interaction strength of the Heisenberg Hamiltonian. Note that in general $\ell$ can change significantly, viz. one order of magnitude. For instance, if the we fix $\ell=100\operatorname{nm}$ and $\lambda\in\mathcal{O}(10^{1})$ nm as before, we find that $\epsilon$ changes one order of magnitude, i.e. $\epsilon\simeq 100$. For materials other than MnSi, e.g. Fe1–xCoxSi or Mn1–xFexGe, $\lambda$ can also be of the order $\mathcal{O}(10^{2})\operatorname{nm}$. Therefore, considering the other two cases where $\lambda$ is fixed as $\lambda\sim 10^{2}\operatorname{nm}$ and $\ell$ takes its maximum and minimum values, we find the relevant ranges for $\epsilon$:
@@
In Secs. 2.5 and 2.6, we saw that in the special physical case of $p=2$, which describes chiral magnets with anisotropy interactions, any spherical magnetic Skyrmion can be represented by the analytical lump solution, regardless of its DM coupling, $\kappa$. However, the potential $V_{p=2}$ corresponds to the sum of a magnetic field and anisotropy (easy-plane) contributions. To realize this potential, the couplings of the anisotropy term must be exactly half of that of the Zeeman term. It turns out to be possible [55]. Furthermore, since the critical coupling regime corresponds to the condition $\forall\leavevmode\nobreak\ \ell,\ \epsilon=2$, which satisfies the inequalities (2.76), we suggest that, in principle, a BPS magnetic Skyrmion could be realized experimentally.
@@ 3 Conclusion
In this paper, we have studied magnetic Skyrmions in chiral magnets with three different potentials, corresponding to $p=1$, $p=\frac{3}{2}$ and $p=2$ in the limits of the control parameter $\epsilon=\frac{\kappa^{2}}{h}=\frac{4D^{2}}{|B|J}$ becoming very large and going to zero. The $p=1$ case is simply the Zeeman potential term and the $p=2$ case is a particular combination of the easy-plane (anisotropy) and the Zeeman potentials. The $p=\frac{3}{2}$ case is probably somewhat more academic. The $\epsilon$ vanishing limit reduces the Skyrmion to a sigma model lump with a lump size that we accurately estimated, except for the case of the Zeeman term ($p=1$). The large-$\epsilon$ limit instead corresponds to the “restricted” magnetic Skyrmions, for which we found the exact analytic solutions as well. In the $p=\frac{3}{2}$ case, the restricted solution is a compacton, similar in nature to those found in the baby-Skyrme [43, 44, 45] and 3D Skyrme models [46]. However, in the $p=1$ case corresponding to the Zeeman term, we find a new type of restricted solution, which we dub a supercompacton, since it becomes discontinuous. We provide numerical solutions showing that they slowly tend to the supercompacton limit for very large values of $\epsilon$. We have also reviewed the beautiful result of the critically coupled magnetic Skyrmion of Barton-Singer-Ross-Schroers [19], where an analytic solution is available, hence providing many studies with a clear advantage. Their solution is known to be an exact solution for any value of the coupling $\epsilon$ [19], but we find here that dropping the boundary term that leads to the correct lump size for this solution also changes the energy from the lump energy for all values of $\ell$ to the energy functional $4\pi(1-\epsilon)$. This energy functional changes sign at $\epsilon=1$ suggesting $\epsilon=1$ being the critical coupling for having the Skyrmionic phase.